# Solving and learning advective multiscale Darcian dynamics with the Neural Basis Method

### v1

@@ Methods
Across all cases, NBM–OL learns a map from a parameterization (and time, when applicable) to the solution coefficients in the neural basis space, enabling rapid inference of pressure, flux/velocity, and concentration fields for new parameter instances.
Common workflow. Given a parameter input $\boldsymbol{\xi}$ (and time $t$ when relevant), the overall procedure is as follows. When POD compression is not used, steps (ii) and (iii) are omitted.
  1. (i)
    define a parametrization of the operator family;
  2. (ii)
    generate solution coefficient snapshots by solving sampled training instances using NBM;
  3. (iii)
    compress the solution coefficient snapshots with POD to define a latent coordinate $\mathbf{z}$;
  4. (iv)
    train an MLP operator $\mathcal{F}$ to predict $\mathbf{z}$ using a self-supervised objective based on residual $\mathcal{E}_{\mathrm{rel}}$;
  5. (v)
    at inference, evaluate $\mathcal{F}$ to obtain $\mathbf{z}$, reconstruct full solution coefficients via POD projection, and recover fields at collocation locations by evaluating the neural basis expansion. When needed, the velocity field is obtained from the reconstructed mass flux and density.
Case-I: permeability-varying Darcy flow operator. We parametrize permeability heterogeneity using a piecewise-constant $5\times 5$ blockwise field (block25). Each block25 realization is generated from a Gaussian random field model (SI), controlled by the heterogeneity level, correlation length(s), and a rotation angle orienting the principal correlation direction. Given a sampled block25 instance, we apply it to the base permeability template $\kappa_{\mathrm{ref}}(\mathbf{x})$ via blockwise multiplication to obtain $\kappa(\mathbf{x};\boldsymbol{\xi}_{\kappa})$. For each training sample, NBM produces steady coefficient snapshots $(\boldsymbol{\theta}_{p},\boldsymbol{\theta}_{u})$; we apply POD to $\boldsymbol{\theta}_{p}$ and $\boldsymbol{\theta}_{u}$ separately and define $\mathbf{z}=[\mathbf{z}_{p};\mathbf{z}_{u}]$. We learn the mapping $\mathcal{F}(\boldsymbol{\xi}_{\kappa})\mapsto\mathbf{z}$ using a two-branch multilayer perceptron (MLP), with one branch predicting $\mathbf{z}_{p}$ and the other predicting $\mathbf{z}_{u}$, trained via a residual-based objective obtained by evaluating the Darcy flow NBM residual in the reduced coordinate system. At inference, we reconstruct the steady pressure and velocity fields from the coefficients predicted by the network.
Case-II: boundary-flux-varying Darcy flow operator. We parametrize the operator family by a scalar control $g^{\mathrm{top}}$ specifying the outward top-boundary mass-flux. For each sampled $g^{\mathrm{top}}_{i}$, we run NBM to generate a coefficient trajectory over the Darcy time horizon, $\{\boldsymbol{\theta}(t^{n};g^{\mathrm{top}}_{i})\}$, where $\boldsymbol{\theta}=[\boldsymbol{\theta}_{\mathrm{div}};\boldsymbol{\theta}_{\mathrm{curl}};\boldsymbol{\theta}_{p}]\in\mathbb{R}^{3N_{b}}$ collects the structured Helmholtz components of mass-flux together with pressure. We apply POD to $\boldsymbol{\theta}_{\mathrm{div}}$, $\boldsymbol{\theta}_{\mathrm{curl}}$, and $\boldsymbol{\theta}_{p}$ separately to obtain $\mathbf{z}=[\mathbf{z}_{\mathrm{div}};\mathbf{z}_{\mathrm{curl}};\mathbf{z}_{p}]$. We learn a time-conditioned latent map $\mathbf{z}(t)=\mathcal{F}(g^{\mathrm{top}},t)$ using three MLP branches (one per latent block), trained by minimizing the Darcy flow NBM residual in the reduced coordinate system. At inference, for new $g^{\mathrm{top}}$ and $t$, we reconstruct $\boldsymbol{\theta}(t)$ and recover pressure and velocity fields at collocation locations.
Case-III: boundary-flux/concentration-varying Darcy–transport operator. We design a coupled Darcy–transport setting with two scalar controls: $g^{\mathrm{top}}$ (top-boundary outward mass-flux) and $c^{\mathrm{left}}$ (left-boundary injected tracer concentration). For each sampled pair $(g^{\mathrm{top}}_{i},c^{\mathrm{left}}_{i})$, we directly use the pretrained Case-II model to infer the velocity sequence $\{\mathbf{u}^{n+1}(g^{\mathrm{top}}_{i})\}$ at coarse Darcy time steps indexed by $n$. Each coarse Darcy step ($\Delta T$) is split into $N_{\mathrm{tr}}$ transport substeps with $\Delta t=\Delta T/N_{\mathrm{tr}}$, and NBM is used to generate concentration coefficient snapshots $\{\boldsymbol{\theta}_{c}(t^{m+1};g^{\mathrm{top}}_{i},c^{\mathrm{left}}_{i})\}$ over all substeps indexed by $m$. We apply POD to $\{\boldsymbol{\theta}_{c}\}$ to define a latent concentration state $\mathbf{z}_{c}$. We then learn a single time-conditioned latent map $\mathbf{z}_{c}(t)=\mathcal{F}(g^{\mathrm{top}},c^{\mathrm{left}},t)$, trained by minimizing the transport NBM residual expressed in latent form. At inference, we obtain the Darcy velocity sequence from the pretrained Case-II model, evaluate $\mathcal{F}$ to reconstruct $\boldsymbol{\theta}_{c}(t)$, and recover concentration fields at collocation locations over the time horizon.

### v2

@@ Appendix H NBM–OL implementation and experiment configuration
The NBM–OL implementation follows the structured workflow summarized in Fig. 4. We consider three operator-learning settings: Case-I, a permeability-varying steady Darcy flow operator; Case-II, a boundary-flux-varying Darcy flow operator; and Case-III, a boundary-flux/concentration-varying Darcy–transport operator. Across all cases, NBM–OL learns a map from the problem parameters (and time, when applicable) to solution coefficients in the neural basis space, enabling rapid inference of pressure, flux/velocity, and concentration fields.
The common workflow is as follows: (i) define a parameterization of the operator family; (ii) generate coefficient snapshots by solving sampled training instances using NBM; (iii) apply POD blockwise to compress the coefficient snapshots into latent coordinates; (iv) train an MLP operator in the latent space using the residual metric $\mathcal{E}_{\mathrm{rel}}$ as a self-supervised objective; and (v) at inference, predict the latent coordinates, reconstruct the full coefficients, and recover the solution fields by neural basis expansion. When POD compression is not used, steps (ii) and (iii) are omitted.
All experiments use the square domain $\Omega=[0,762]~\mathrm{m}\times[0,762]~\mathrm{m}$. Interior residuals are evaluated on a structured $100\times 100$ collocation set (10,00010{,}000 points), and boundary residuals on 400 uniformly distributed boundary points with corners excluded. Pressure, mass-flux (or velocity in the incompressible case), and concentration are represented by separate neural bases, all with size $N_{b}=800$. When POD is used, it is applied separately to each coefficient block. OOD tests are constructed by extending one or more parameter axes and/or the time horizon beyond the training ranges.
Case-I: permeability-varying steady Darcy flow operator. We impose mixed boundary conditions: $p=69~\mathrm{bar}$ on $x=0$, $u_{x}=0.3~\mathrm{m~day^{-1}}$ on $x=762~\mathrm{m}$, and $u_{y}=1.5~\mathrm{m~day^{-1}}$ on $y=0$ and $y=762~\mathrm{m}$, using the outward-normal sign convention. The base permeability template $\kappa_{\mathrm{ref}}(\mathbf{x})$ is defined on a $100\times 100$ grid, and heterogeneity is parameterized by a piecewise-constant $5\times 5$ blockwise field (block25) generated from a Gaussian random field. Each block25 realization is upsampled by block repetition and applied multiplicatively to $\kappa_{\mathrm{ref}}(\mathbf{x})$. We generate $N_{s}=1000$ realizations and compute steady NBM coefficient snapshots $(\boldsymbol{\theta}_{\mathrm{div}},\boldsymbol{\theta}_{p})$; only the divergence-free velocity block is needed because this setup is incompressible. POD retains $r_{p}=16$ and $r_{q}=16$ modes, yielding $\mathbf{z}=[\mathbf{z}_{p};\mathbf{z}_{q}]\in\mathbb{R}^{32}$. The latent map is learned by a two-branch MLP, one branch for pressure and one for velocity, each with two hidden layers of width 64 and $\tanh$ activation. Training uses Adam with learning rate $10^{-3}$, batch size 16, and 500 epochs. OOD tests are generated by perturbing the Gaussian random field controls, including variance/contrast scaling, correlation-length changes, and rotation-angle shifts.
Case-II: boundary-flux-varying Darcy flow operator. The operator family is parameterized by a scalar top-boundary mass-flux control $g_{\mathrm{top}}$, uniformly sampled from $488,976]~\mathrm{kg~day^{-1}~m^{-2}}$. We prescribe $p=103~\mathrm{bar}$ on $x=0$ and impose no-flux conditions on the right and bottom boundaries. The permeability field is fixed and multiscale on a $100\times 100$ grid. We generate $N_{s}=200$ scenarios, each simulated for $N_{\mathrm{Darcy}}=20$ Darcy time steps with $\Delta T=10.0~\mathrm{days}$, yielding $N_{\mathrm{snap}}=4000$ snapshots for each coefficient block $(\boldsymbol{\theta}_{\mathrm{div}},\boldsymbol{\theta}_{\mathrm{curl}},\boldsymbol{\theta}_{p})$. POD retains $r_{\mathrm{div}}=8$, $r_{\mathrm{curl}}=8$, and $r_{p}=8$ modes, giving $\mathbf{z}=[\mathbf{z}_{\mathrm{div}};\mathbf{z}_{\mathrm{curl}};\mathbf{z}_{p}]\in\mathbb{R}^{24}$. The latent operator $\mathcal{F}(g_{\mathrm{top}},t)$ is represented by three MLP branches, each with two hidden layers of width 32 and $\tanh$ activation. Training uses Adam with learning rate $10^{-3}$, batch size 16, and 500 epochs. OOD evaluation extends the $g_{\mathrm{top}}$ range by $50\%$ and the prediction horizon by $100\%$ beyond the training ranges.
Case-III: boundary-flux/concentration-varying Darcy–transport operator. The coupled operator family is parameterized by $(g_{\mathrm{top}},c_{\mathrm{left}})$, where $g_{\mathrm{top}}$ is the top-boundary Darcy mass-flux control and $c_{\mathrm{left}}$ is the injected tracer concentration on the left boundary, sampled from $\mathcal{U}(50.0,100.0)~\mathrm{ppm}$. The Darcy boundary setup and fixed permeability field are the same as in Case-II unless stated otherwise. We generate $N_{s}=80$ independent scenarios. For each $g_{\mathrm{top}}$, the pretrained Case-II model is used to infer the Darcy velocity sequence, and each Darcy time step is split into $N_{\mathrm{tr}}=8$ transport substeps. We use $N_{\mathrm{Darcy}}=10$ Darcy coarse steps, so the total number of transport substeps is $N_{\mathrm{sub}}=N_{\mathrm{Darcy}}N_{\mathrm{tr}}$. POD is applied to the transport coefficient snapshots $\{\boldsymbol{\theta}_{c}\}$ and retains $r_{c}=20$ modes, yielding $\mathbf{z}_{c}\in\mathbb{R}^{20}$. The latent map $\mathbf{z}_{c}(t)=\mathcal{F}(g_{\mathrm{top}},c_{\mathrm{left}},t_{\mathrm{norm}})$ is learned by a single MLP with two hidden layers of width 64, $\tanh$ activation, and weight normalization. Training uses Adam with learning rate $5\times 10^{-3}$, batch size 32, and 1000 epochs. OOD evaluation extends both $(g_{\mathrm{top}},c_{\mathrm{left}})$ by $50\%$ and the prediction horizon by $100\%$ beyond the training ranges.
Appendix I Physics-informed neural network (PINN) baseline
We use a two-network, two-stage vanilla PINN baseline for the coupled Darcy–transport problem. In Stage A, pressure is approximated by a neural network $p(\cdot;\omega_{p})$ and trained by minimizing a weighted sum of mean-squared residuals at collocation points, consisting of the strong-form slightly compressible Darcy residual in the interior together with boundary penalties enforcing the left Dirichlet pressure condition and mass-flux conditions on the other boundaries, where the normal mass flux is obtained from $\nabla p$ via automatic differentiation. In Stage B, concentration is approximated by a second network $c(\cdot;\omega_{c})$; after freezing $\omega_{p}$, we train $\omega_{c}$ using a conservative transport residual at space–time collocation points, together with penalties enforcing the initial condition $c(\cdot,0)=0$ and the time-dependent inlet Dirichlet condition on the left boundary. To reduce cost while preserving PDE-level coupling, once Stage A converges we precompute the mass-flux components and their divergence induced by $p(\cdot;\omega_{p})$ on the full space–time collocation set and reuse them throughout Stage B, so that only $c(\cdot;\omega_{c})$ is differentiated during transport training.
Both networks are fully connected MLPs with $\tanh$ activations, width $64$, depth $4$, and scalar output, taking $(x,y,t)$ as input. Collocation uses a fixed $50\times 50$ spatial grid of cell centers and $N_{t}=10$ time levels, giving $50\times 50\times 10$ interior space–time points, while boundary collocation points are the boundary cell centers replicated across the same time levels. For Stage A, we use Adam with learning rate $10^{-3}$ for 20,00020{,}000 iterations with mini-batches of size $4096$ for interior points and $1024$ for each boundary set, followed by full-batch L-BFGS for 2,0002{,}000 iterations using strong Wolfe line search, history size $30$, $\texttt{tolerance\_grad}=10^{-10}$, and $\texttt{tolerance\_change}=10^{-12}$; the loss weights are $(w_{\mathrm{pde}},w_{L},w_{R},w_{B},w_{T})=(1,1,1,1,10)$. For Stage B, we use Adam with learning rate $10^{-3}$ for 10,00010{,}000 iterations with mini-batches of size $4096$ for interior points, $1024$ for the initial condition, and $1024$ for the inlet condition, followed by full-batch L-BFGS for 2,0002{,}000 iterations with the same settings; the loss weights are $(w_{\mathrm{pde}}^{(c)},w_{\mathrm{ic}},w_{\mathrm{in}})=(1,0.01,10)$. For the multiscale cases, both Stage A and Stage B use 50,00050{,}000 Adam iterations followed by 20,00020{,}000 L-BFGS iterations.
CRediT author statement
Yuhe Wang: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Visualization, Writing - Original Draft, Review & Editing. Min Wang: Conceptualization, Methodoloty, Formal analysis, Investigation, Writing - Review & Editing.